OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10 Answer choices: 2 -1 1 -2 Thanks!

OpenStudy (abb0t):

Well, you said you tried, so...my next question to you is, what are \(your\) answers?

OpenStudy (anonymous):

The answers i came up with arent any of the choices .

OpenStudy (abb0t):

For any of them? Lol right... Anyways, what method are you using? Are you using elimination, substitution, or are you doing this using matrices; cramers rule?

OpenStudy (anonymous):

Substitution

OpenStudy (abb0t):

I'll help you with the first one, but you will have to do the others on your own. I am not going to do your quiz or homework for you.

OpenStudy (abb0t):

I think for this, you might want to use elimination method to avoid use of fractions. So, you have a system: \(\left\{\begin{matrix} 2x + 3y = 1 & & \\ 3x-4y=10& & \end{matrix}\right.\) \(\Rightarrow\) \(\left\{\begin{matrix}4 (2x + 3y = 1) & & \\ 3 (3x-4y=10)& & \end{matrix}\right.\) basically you're multiplying to eliminate one term, in this case, "y" so you can get "x" only.

OpenStudy (anonymous):

okay

OpenStudy (abb0t):

now, you get \(\Rightarrow\) \(\begin{matrix} 8x+12y=4 & & \\ 9x-12y=30 & & \end{matrix}\)

OpenStudy (abb0t):

Notice that the "y" term cancels out, and you're left with: \(17x=34\) solving for "x", your answer is \(\large 2\)

OpenStudy (anonymous):

Ohh ok thank you .